OpenStudy (anonymous):

Evaluate the integral (−43x) dx

zepdrix (zepdrix):

\[\large \int\limits -43x\;dx \qquad = \qquad -43\int\limits x\;dx\] We can pull the constant coefficient outside of the integral, it won't affect the process at all. From here, use the `Power Rule for Integrals`\[\large \int\limits x^n\;dx \qquad = \qquad \frac{x^{n+1}}{n+1}+C\] ~Increase the power by 1 ~Divide by this new power.

zepdrix (zepdrix):

So it appears you have x^1 Understand how to proceed? :o

OpenStudy (anonymous):

if I have x^1 would it just be 0?

zepdrix (zepdrix):

No. With integrals, the power INCREASES, not decreases. \[\large \int\limits x^1 \;dx \qquad = \qquad \frac{x^{1+1}}{1+1}\]Which we can write this way,\[\large \frac{1}{2}x^2\]

OpenStudy (anonymous):

oh ok I understand now. I see what I did wrong. Thank you